OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -14, and 5 + 8i f(x) = x4 - 362.5x2 + 1450x - 4984 f(x) = x4 - 9x3 + 32x2 - 725x + 4984 f(x) = x4 - 67x2 + 1450x - 4984 f(x) = x4 - 9x3 - 32x2 + 725x - 4984

OpenStudy (anonymous):

\[x-4)(x+14)(quadratic)\] where 'quadratic' has zeros of \(5+8i\) and \(5-8i\) which is the only hard part, and it is not that hard to find

OpenStudy (anonymous):

you have three choices. one is hard, one is easy and one is really really easy you pick

OpenStudy (anonymous):

I'm really confused still, what?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

you are given three zeros right?

OpenStudy (anonymous):

4, -14, and 5 + 8i

OpenStudy (anonymous):

in fact you are actually given 4 zeros, because if \(5+8i\) is a zero, then so is its conjugate \(5-8i\)

OpenStudy (anonymous):

since 4 is a zero, one factor of the polynomial is \(x-4\)

OpenStudy (anonymous):

since \(-14\) is a zero, one factor of the polynomial is \(x-(-14) =x+14\)

OpenStudy (anonymous):

clear so far or no?

OpenStudy (anonymous):

@satellite73 yeah i get that!

OpenStudy (anonymous):

then the next part is to find a quadratic with zeros of \(5+8i\) and \(5-8i\)

OpenStudy (anonymous):

@satellite73 x^2-10x-80i+89???

OpenStudy (anonymous):

whoa hold the phone there should be no \(i\) in your answer

OpenStudy (anonymous):

you have a choice one way is to work backwards \[x=5+8i\\ x-5=8i\] then square (carefully) and get \[(x-5)^2=(8i)^2\\ x^2-10x+25=-16\]

OpenStudy (anonymous):

and so your quadratic is \[x^2-10x+41\]

OpenStudy (anonymous):

oh ok @satellite73

OpenStudy (anonymous):

then multiply all that mess out \[(x+4)(x-14)(x^2-10x+41)\]

OpenStudy (anonymous):

@satellite73 I got x^4-10x^3-25x^2+100x-2146?

OpenStudy (anonymous):

i would use this so because otherwise i am sure i would make an algebra mistake http://www.wolframalpha.com/input/?i=%28x%2B4%29%28x-14%29%28x^2-10x%2B41%29 look under "expanded form"

OpenStudy (anonymous):

oh crap i made a mistake!!

OpenStudy (anonymous):

guess i don't know how to square myself \[(x-5)^2=(8i)^2\\ x^2-10x+25=-16\] is wrong \[(x-5)^2=(8i)^2\\ x^2-10x+25=64\] is more like it

OpenStudy (anonymous):

yeah i was wondering because that's not an answer choice

OpenStudy (anonymous):

so \[x^2-10x+89\] is your polynomial

OpenStudy (anonymous):

for which one? @satellite73

OpenStudy (anonymous):

the one with the two complex zeros now go to \[(x+4)(x-14)(x^2-10x+89)\]

OpenStudy (anonymous):

btw if you ever have to do another one of these, and are told that a zero is \(a+bi\) then the quadratic is \[x^2-2ax+(a^2+b^2)\]

OpenStudy (anonymous):

I'm confused because that's still none of my answer choices @satellite73

OpenStudy (anonymous):

we had \(5+8i\) and the quadratic was \[x^2-2\times 5x+(5^2+8^2)\\ x^2-10x+25+64\\ x^2-10x+89\]

OpenStudy (anonymous):

yeah because i am really an idiot today i wrote "if 4 is a zero then \(x-4\) is a factor, but then wrote the factor as \(x+4\)which is wrong lets try one more time\[(x-4)(x+14)(x^2-10x+89)\]

OpenStudy (anonymous):

That's more like it! Thank you soooo much for taking the time to help me!(:

OpenStudy (anonymous):

i sure as hell hope \[x^4-20 x^3+133 x^2-330 x-4984\]is an answer choice!

OpenStudy (anonymous):

yw sorry i screwed up twice, but i hope you get the idea